# Feature Representation Transferring to Lightweight Models via Perception Coherence

### v1

@@ Appendix E Experiment details
$d(u,v)=\frac{1}{2}\left(1-\frac{\langle u,v\rangle}{\|u\|\cdot\|v\|}\right)\ .$ (E.1)
@@ E.1 Toy dataset on 2D and 3D
For these experiments, we use scikit-learn package in Python to simulate the datasets. The two-moon dataset has 700 points, and the 5 $2D$ Gaussian clusters have 1000 points (Fig. E.1). The 3D clusters also have 1000 points. The student configuration is randomly initialized using Gaussian distribution. For training, we use batch size equal to 64, and learning rate equal to 0.1. We use Adam optimizer of the Pytorch package for training during 800 epochs. We set the temperature equal to $0.1$ in the soft ranking function.
Figure E.1: Matching process on $2D$ Gaussian datasets. We observe that learned configuration preserve a global structural coherence without preserving completely the geometry of the teacher configuration.
@@ E.2 Retrieval experiments on CIFAR10 and CUB-200
Following [22], we use ResNet18 as the teacher model for both datasets. We also use the same student models for each datasets, shown in Fig. E.2. We apply our method only on the penultimate layer (dim 64 in ig. E.2), without using the final classification layer.
@@
(a) Student model for CIFAR10.
(b) Student model for CUB200.
Figure E.2: Student models for CIFAR10 and CUB200. Images are extracted from [22]. Notice that we perform transferring on the features of penultimate layer (dimension of 64), and do not use the final classification layer.
@@ E.3 Classification on CIFAR-100
For teacher models, we use pretrained models released in the official code of the work [29] without any modification. We also use the same student models from the same code. We follow the same training scheme as the work [29]. More precisely, for optimizer, we use SGD(lr=0.01, momentum=0.9, weight_decay=5e-4). We use the training set for transferring, during 240 epochs, where we decay the learning by a factor of 0.1 at epochs 150, 180 and 210. The batch size is 64. Data augmentation techniques include random cropping (32x32, padding=4) and random horizontal flipping. We fix temperature $\tau=0.2$ for teacher model and $\tau=0.3$ for student models. Our technique is applied on the penultimate layer and the logit layer (before softmax function). The total loss for training the classification student model is $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cross-entropy}}+\lambda\mathcal{L}_{\text{ours}}$, where we fix $\lambda=5$. We perform 5 independent runs to report results in Table 4.3. The running time for one run (240 epochs) is as follows: ShuffleNetV1: about 45 minutes. ShuffleNetV2: about 52 minutes. MobileNetV2: about 1 hour.
E.4 Ablation study: effect of batch size through qualitative evaluation
We use the toy datasets where we transfer from $3D\to 2D$ (Fig. E.3) and $2D\to 2D$ (Fig. E.4). We use different batch sizes ($B$) to observe the effect of batch size on the learned configuration. We see that when $B$ is too small ($B=3$), the learned configuration does not preserve a structural coherence of the reference one. This is to be expected, as a too small $B$ provides very little information about relation between data points. From $B=7$, the learned configuration preserves more or less the reference one. The results start to stabilize from $B=15$. It is reasonable to conjecture that we do not need too large $B$ (compared to the dataset size) to have the learned configuration preserving the coherence of the reference one. This is because in each training epoch, we randomly sample in each batch $B$ points, until finishing to process the whole dataset. Hence, each batch consider the relation between each point (reference point) to other $B-1$ points (compared set of points). Due to randomness of sampling, in each epoch, each point is compared to a different compared set of $B-1$ points. Hence, after different epochs, we should somehow have the global structural information between data points, without using a too large batch-size.
(a) Reference configuration.
(b) Batch size equal to 3.
(c) Batch size equal to 7.
(d) Batch size equal to 15.
(e) Batch size equal to 32.
(f) Batch size equal to 64.
(g) Batch size equal to 128.
Figure E.3: Ablation study of the effect of batch size on $3D\to 2D$ gaussian dataset, where we train with different batch sizes.
(a) Batch size equal to 3.
(b) Batch size equal to 7.
(c) Batch size equal to 15.
(d) Batch size equal to 32.
(e) Batch size equal to 64.
(f) Batch size equal to 128.
Figure E.4: Ablation study of the effect of batch size on two-moon dataset, where we train with different batch sizes.

### v2

@@ Appendix E Experiment details
$d(u,v)=\frac{1}{2}\left(1-\frac{\langle u,v\rangle}{\|u\|\cdot\|v\|}\right)\ .$ (12)
@@ E.1 Toy dataset on 2D and 3D
For these experiments, we use scikit-learn package in Python to simulate the datasets. The two-moon dataset has 700 points, and the 5 $2D$ Gaussian clusters have 1000 points. The 3D clusters also have 1000 points. The student configuration is randomly initialized using Gaussian distribution. For training, we use batch size equal to 64, and learning rate equal to 0.1. We use Adam optimizer of the Pytorch package for training during 800 epochs. We set the temperature equal to $0.1$ in the soft ranking function.
@@ E.2 Retrieval experiments on CIFAR10 and CUB-200
Following Passalis et al. (2020a), we use ResNet18 as the teacher model for both datasets. We also use the same student models for each datasets, shown in Fig. 9. We apply our method only on the penultimate layer (dim 64 in Fig. 9), without using the final classification layer.
@@
Figure 9: Student models for CIFAR10 and CUB200. Images are extracted from Passalis et al. (2020a). Notice that we perform transferring on the features of penultimate layer (dimension of 64), and do not use the final classification layer.
@@ E.3 Classification on CIFAR-100
For teacher models, we use pretrained models released in the official code of the work Tian et al. (2020) without any modification. We also use the same student models from the same code. We follow the same training scheme as the work Tian et al. (2020). More precisely, for optimizer, we use SGD(lr=0.01, momentum=0.9, weight_decay=5e-4). We use the training set for transferring, during 240 epochs, where we decay the learning by a factor of 0.1 at epochs 150, 180 and 210. The batch size is 64. Data augmentation techniques include random cropping (32x32, padding=4) and random horizontal flipping. We fix temperature $\tau=0.2$ for teacher model and $\tau=0.3$ for student models. Our technique is applied on the penultimate layer and the logit layer (before softmax function). The total loss for training the classification student model is $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cross-entropy}}+\lambda\mathcal{L}_{\text{ours}}$, where we fix $\lambda=5$. We perform 5 independent runs to report results in Table 3. The running time for one run (240 epochs) is as follows: ShuffleNetV1: about 45 minutes. ShuffleNetV2: about 52 minutes. MobileNetV2: about 1 hour.
E.4 Ablation study: effect of scaling temperature in the soft ranking
In this section, we investigate the effect of scaling temperature in the soft ranking. We fix the temperature of the teach models as precised in Section E.2, and use different temperatures for the student model. Results are shown in Table 6.
Table 6: Ablation study: performance model with different values of $\tau$.
$\tau$ (temperature) 0.001 0.01 0.05 0.1 0.3 0.5 5.0 10.0
CUB200 subset mAP 15.61 16.93 14.30 17.24 28.42 30.38 14.16 13.17
top-10 22.67 24.29 20.11 23.86 36.55 38.49 18.87 18.04
CIFAR10 mAP 50.22 49.36 49.50 50.86 54.25 53.72 27.38 25.14
top-100 61.02 60.27 60.22 61.36 65.00 63.89 32.43 28.51
From Table 6, we have the following remarks:
  • Small Temperature. With very small $\tau$, the sigmoid approaches a step function, causing vanishing gradients almost everywhere. This was confirmed in our ablation study on CUB200, where using $\tau\in\{0.001,0.01,0.05\}$ consistently led to degraded performance.
  • Large Temperature. Conversely, large $\tau$ (e.g., $\tau\geq 5$) causes the sigmoid $\Lambda$ to behave nearly linearly, with: $(x)\approx\frac{1}{2}+\frac{x}{4\tau}\quad\text{for small }|x/\tau|$. This weakens the ranking effect and makes the method rely more on raw dissimilarity magnitudes, which is suboptimal in lightweight settings (e.g., when $f_{1}\ll f_{2}$ in capacity). We observed performance drops in both CIFAR100 and CUB200 in this case.
  • Summary. In practice, intermediate values $\tau\in\{0.1,0.3,0.5\}$ yield the best results. This confirms the need for careful temperature tuning.